OpenStudy (anonymous):

Find another name for the given number but with the denominator indicated. Use multiplication by 1. 8/9=?/45

hero (hero):

\[\frac{8}{9} = \frac{x}{45}\] Cross multiply: \[(9)x = 8(45)\] Divide both sides by 9: \[x = \frac{8(45)}{9}\]\[x = 8(5)\]\[x = 40\]